OpenStudy (anonymous):

Let f(x)=(x−5)2. Find the average rate of change of f(x) with respect to x from x=a to x=a+h if h≠0. Simplify your answer as much as possible.

OpenStudy (anonymous):

i am going to guess this is \[f(x)=(x-5)^2=x^2-10x+25\]